OpenStudy (anonymous):

if f(x) = −4x − 1, find f(−2). −9 −7 −10 7

OpenStudy (anonymous):

need help n explain it to me thxs

OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2). y= -4(-2)-1 y= 8-1 y= 7

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

ur welcome

OpenStudy (radar):

x=-2 f(-2)=-4(-2) - 1 f(-2)=8-1 f(-2)=7

OpenStudy (anonymous):

thank you sir

OpenStudy (radar):

You are welcome, HeyltaAri16 beat me to the draw lol

OpenStudy (anonymous):

hahaha lol :D